I need help on the following question. I don't understand it. (5-y)
+(6-y)-(5-y)=0
Help

Answer by checkley71(8403)

(5-Y)+(6-Y)-(5-Y)=0
5-Y+6-Y-5+Y=0
-Y=5-5-6
-Y=-6
Y=6 ANSWER.
PROOF
(5-6)+(6-6)-(5-6)=0
-1+0+1=0
0=0